OpenStudy (anonymous):

Multiply and then simplify the radical. 5 square root of 10 times 3 square root of 8 A. 15 square root of 80 B. 15 square root of 18 C. 45 square root of 5 D. 60 square root of 5

OpenStudy (anonymous):

A

jhonyy9 (jhonyy9):

not is right D will be because we get 15sqrt80 = 15sqrt16*5 =15*4sqrt5=60sqrt5

jhonyy9 (jhonyy9):

so A

jhonyy9 (jhonyy9):

and D are right too

OpenStudy (anonymous):

=)